a piece of 12 ft wire is folded into a rectangle that is 3 ft longer than it is wide. what’s the area?
Answer by Boreal(15235)

perimeter=12ft.
half-perimeter=6 ft= L+W, and L=W+3
so W+3+W=6
2W=3
W=1.5 feet
L=4.5 feet
A=their product or 6.75 ft^2.